OpenStudy (anonymous):

“To pass by me you must tell me how to convert standard form into the general, vertex form... I have a test on it next week.”

OpenStudy (anonymous):

Explain how to convert f(x) into the general, vertex form of the equation. Use complete sentences. You may use the f(x) you created in question 1 as an example.

OpenStudy (anonymous):

my f(x) = x^2 + 10x + 16

hartnn (hartnn):

do you know how to complete the square ?

OpenStudy (anonymous):

That's with using the quadratic formula?

hartnn (hartnn):

no... its like squaring the half of co-efficient of x

OpenStudy (anonymous):

Ok, How would I do that?

hartnn (hartnn):

f(x) = x^2 + 10x + 16 whats the co-efficient of x ?

OpenStudy (anonymous):

10? Because it's before x?

hartnn (hartnn):

yes

hartnn (hartnn):

now divide it by 2 and square it what u get ?

OpenStudy (anonymous):

25?

hartnn (hartnn):

yes so add and subtract 25 on right side (x^2+10x +25) + 16-25 did u get what i did ?q

OpenStudy (anonymous):

ok... (x^2 + 10x +25) - 9 I don't think I did that right, do I remove the parenthesis?

hartnn (hartnn):

i kept the paranthesis so that it becomes easy to identify that it is a perfect square trinomial

hartnn (hartnn):

y = (x+5)^2 - 9

OpenStudy (anonymous):

Oh ok, I see where you got that, and this is already the vertex form? Do I have to put the 'a' variable on the outside?

hartnn (hartnn):

if you need to find 'a' then put 1 outside, because a comes out to be =1 y= a (x-h)^2 +k y= (x+5)^2 -9 a=1 h= -5 k = -9

hartnn (hartnn):

got it ? y= (x+5)^2-9 is your vertex form :)

OpenStudy (anonymous):

Oh! Ok, cool, I understand that!:) Thanks so much! By the way, do you also know with standard form, how I would find the solutions for that? Just a quick question.

hartnn (hartnn):

whats the standard form ?

OpenStudy (anonymous):

ax^2 + bx + c

hartnn (hartnn):

f(x) = x^2 + 10x + 16 is already in that form a= 1 b=10 c=16

OpenStudy (anonymous):

Yes, but would there be a way to find the 'solutions' for that equation?

hartnn (hartnn):

oh, you need solution for f(x) ?? [but the question didn't ask for it....] so f(x) = 0 x^2+10x +16 =0 know factor method ?

hartnn (hartnn):

not actually can u tell me 2 numbers whose sum is 10 and product = 16

OpenStudy (anonymous):

2 and 8?

hartnn (hartnn):

yes so x^2+8x +2x +16 =0 know how to proceed ? factor

OpenStudy (anonymous):

No, I'm not sure, this would be factoring it out into two sets of parenthesis, right?

OpenStudy (anonymous):

Something like (x+8) (x + 2)?

hartnn (hartnn):

its like this x (x+8) +2 (x+8) = 0 (x+8) (x+2) = 0 x+8 =0 or x+2 = 0 x= -8 or x= -2 thats it!

hartnn (hartnn):

yes, you were correct :)

OpenStudy (anonymous):

Oh, cool! Thanks! You've really helped! And thanks for the medal:)

hartnn (hartnn):

welcome ^_^